- Find L-1{f(t)} f(t) s+2 s-(s + 1)2
Answer #1

L^{-1}\left\{\frac{s+2}{s^2\left(s+1\right)^2}\right\}\\\mathrm{Take\:the\:partial\:fraction\:of}\:\frac{s+2}{s^2\left(s+1\right)^2}:\quad -\frac{3}{s}+\frac{2}{s^2}+\frac{3}{s+1}+\frac{1}{\left(s+1\right)^2}\\=L^{-1}\left\{-\frac{3}{s}+\frac{2}{s^2}+\frac{3}{s+1}+\frac{1}{\left(s+1\right)^2}\right\}\\\\\mathrm{Use\:the\:linearity\:property\:of\:Inverse\:Laplace\:Transform:}\\\\\mathrm{For\:functions\:}f\left(s\right),\:g\left(s\right)\mathrm{\:and\:constants\:}a,\:b:\quad L^{-1}\left\{a\cdot f\left(s\right)+b\cdot g\left(s\right)\right\}=a\cdot L^{-1}\left\{f\left(s\right)\right\}+b\cdot L^{-1}\left\{g\left(s\right)\right\}\\\\=-L^{-1}\left\{\frac{3}{s}\right\}+L^{-1}\left\{\frac{2}{s^2}\right\}+L^{-1}\left\{\frac{3}{s+1}\right\}+L^{-1}\left\{\frac{1}{\left(s+1\right)^2}\right\}\\L^{-1}\left\{\frac{3}{s}\right\}:\quad 3\\L^{-1}\left\{\frac{2}{s^2}\right\}:\quad 2t L-`1{1/s+a}=e-at , L-`1{1/(s+a)2}=te-at

L^{-1}\left\{\frac{3}{s+1}\right\}:\quad 3e^{-t}\\L^{-1}\left\{\frac{1}{\left(s+1\right)^2}\right\}:\quad e^{-t}t\\\\=-3+2t+3e^{-t}+e^{-t}tanswer